OpenStudy (anonymous):

How to find the equation of the tangent line at (pi,0) for this equation

OpenStudy (anonymous):

\[\frac{ 3x^2+ysinx-3x^2\sin5y }{ 2y+5x^3\cos5y-2y }\]

OpenStudy (anonymous):

thats the derivative I got. I got 0 as the slope

OpenStudy (phi):

if that is the derivative, the y's and sin(y) are all 0 and cos(y)=1 you get 3x^2/5x^3 = 3/(5x) = 3/(5 pi)

OpenStudy (anonymous):

ok, I'm having a little problem with something, when I have something like 3x^2 sin 5x, would I do the sin 5x first then the 3x^2 or would I do it all together

OpenStudy (phi):

exactly what is the question ?

OpenStudy (anonymous):

\[3x^2\sin5x\]

OpenStudy (anonymous):

if I have something like that

OpenStudy (phi):

that is an expression.... what do you want to do with it ?

OpenStudy (anonymous):

lets say I want to know what y is when x is equal to 3, would I do the sin5x first then multiply by the 3x^2 or will I do it all together

OpenStudy (phi):

are you saying \[ y = 3 x^2 \sin 5x \] now evaluate at x = \( \pi\) ?

OpenStudy (anonymous):

yes, bu would I just put it all into my cal or would I do 5 multiplied by pi, then take the sin of that then multiply by 3x^2

OpenStudy (phi):

as you know, you can change the order of multiplication so you will get the same answer if you did sin 5x * 3 *x^2 or 3* sin 5x * x^2 or 3*x^2 * sin 5x

OpenStudy (anonymous):

so It doesn't matter what order I do them in

OpenStudy (phi):

no, but if x is pi , sin(5 pi) = 0 so you don't have to do any more work... the product will be 0

OpenStudy (anonymous):

ok thx

OpenStudy (phi):

Out of curiosity, what is the original equation that you took the derivative of ?